# Improving the Efficiency of Self-Supervised Adversarial Training through Latent Clustering-Based Selection

### v1

@@ 5.3. Computational Benefits
As suggested by the runtime comparisons depicted in Table 1, utilizing smaller, carefully selected unlabeled data can significantly reduce the overall computation costs of SSAT while maintaining competitive, robust accuracy. This section further visualizes the learning curves of SSAT and discusses the computational benefits associated with our proposed methods.
@@
Figure LABEL:fig:convergence depicts the learning curves of vanilla AT and SSAT algorithms on CIFAR-10 with varying ratios of unlabeled data selected from TinyImages. In our setup, one training epoch corresponds to processing $50$K data points, with $10$ steps for PGD training and $20$ for PGD testing. When no additional data is utilized, as shown in Figure LABEL:fig:0per, the highest robust test accuracy is achieved after approximately $75$ epochs. As illustrated in Figures LABEL:fig:10percrandom and LABEL:fig:10perc, incorporating $10\%$ additional data, whether through random selection or curated approaches, extends the optimal convergence point to approximately $100$ epochs. In contrast, as depicted in Figure LABEL:fig:100perc, leveraging $100\%$ of the unlabeled data delays the achievement of peak accuracy to around $185$ epochs. These findings suggest that including extra unlabeled data in SSAT increases computational demands, as evidenced by the longer training convergence time. Strategic selection reduces redundancy in the dataset, allowing the model to achieve optimal performance with fewer training epochs.
@@
Specifically, we observe that the size of the selected unlabeled dataset plays a critical role in determining the time required for the model to achieve peak accuracy. Larger unlabeled datasets usually extend the convergence times of SSAT, whereas smaller datasets facilitate faster convergence, allowing the model to reach optimal accuracy earlier in training. However, continued training beyond this point often leads to robust overfitting. Early stopping offers a computational advantage in scenarios with limited datasets by preventing overfitting with reduced training durations. Nevertheless, early stopping proves less effective in SSAT when working with large-scale unlabeled datasets, as longer training is necessary to achieve peak performance. To mitigate these challenges, selecting a smaller yet strategically curated subset of data significantly reduces the dataset size, enabling faster convergence and minimizing overall training time. These findings underscore the computational benefits of our proposed data selection schemes, providing a practical solution to the challenges of SSAT with extensive datasets.
5.4. Application to Real-World Medical Data
We extend our study to a real-world healthcare application with COVID-19-related datasets. Specifically, the COVIDGR dataset (Tabik et al., 2020) is utilized as the labeled dataset, and the CoronaHack dataset (Govi, 2020) serves as the source of unlabeled data. This extension allowed us to validate the efficacy of the proposed methods in a domain where data scarcity and model robustness are critical considerations.
Setting. The experimental setup consists of multiple stages, leveraging both datasets in conjunction with ResNet-18 as the base model. The COVIDGR dataset contains $852$ chest X-ray images evenly distributed across two classes: $426$ COVID-19 positive cases (confirmed via RT-PCR testing) and $426$ negative cases. This dataset was split into an $80\%$ training set and a $20\%$ test set for evaluation purposes. The CoronaHack dataset, which comprises $5,910$ chest X-ray images, is divided into an $89\%$ training set and an $11\%$ test set. This dataset has an imbalanced class distribution, with $73\%$ of the images classified as COVID-19 positive and $27\%$ as negative (non-COVID). For our experiments, we strategically select a subset of the training data from the CoronaHack dataset as unlabeled data to assess the effectiveness of our adversarial robustness techniques. We applied the LCS-KM selection strategy to identify and incorporate a minimal yet representative subset of unlabeled data from the CoronaHack dataset into the training process. This approach enabled us to maintain robust model performance while significantly reducing the volume of external data required. The models are trained on SSAT with varying unlabeled data selection schemes using a ResNet-18 architecture. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$ and vanilla PGD-based adversarial training.
@@ 5.4. Application to Real-World Medical Data
Result Analysis. The results of our experiments confirm the effectiveness of strategic data selection in enhancing robust accuracy and optimizing training efficiency. For instance, using LCS-KM to select $10\%$ of the unlabeled data achieved a robust accuracy of $56\%$, compared to $53\%$ when the same proportion was selected randomly, as shown in Figures LABEL:fig:10perrandom and LABEL:fig:10per. Additionally, smaller, strategically selected datasets led to faster convergence, underscoring the benefits of early stopping. Specifically, the best accuracy of $56\%$ was reached at epoch $36$ when using $10\%$ of the data selected with LCS-KM in Figure LABEL:fig:10per. Figure LABEL:fig:20per shows that incorporating $20\%$ of the data with LCS-KM improved the accuracy to $57\%$ at epoch $47$. In contrast, training on the entire unlabeled dataset (i.e., $\alpha=100\%$) required $60$ epochs to achieve the highest robust accuracy of $58\%$ depicted in Figure LABEL:fig:100per. These findings demonstrate that smaller but strategically selected subsets of data can achieve robust accuracy comparable to that of the full dataset while significantly reducing computational time and memory requirements. The success of our approach in this real-world application highlights its practical relevance, particularly in medical diagnostics, where labeled data is limited, yet robust model performance is essential.
6. Additional Analyses
In this section, we conduct additional analyses and experiments to assess the generalizability of our proposed method. We start by visualizing and comparing the latent-space behavior of PCS, LCS-KM, and LCS-GMM to evaluate their effectiveness in selecting critical data points (Section 6.1). Next, we explore the influence of key ratio hyperparameters in our data selection schemes on the robustness performance achieved by SSAT (Section 6.2). Additionally, we evaluate the flexibility of our method across various adversarial training algorithms and perturbation configurations (Section 6.3). Finally, we examine the impact of intermediate model quality, particularly under the regime with insufficient labeled data (Section 6.4). Without explicit mention, we use our LCS-KM method for selecting critical unlabeled data, as it derives the best-performing model when integrated with SSAT as suggested by Table 1.
6.1. Comparative Visualizations
To facilitate a comprehensive understanding of the proposed data selection schemes, we perform a comparative visualization using the TinyImages dataset in the latent space. The latent representations, extracted from the penultimate layer of a pre-trained WRN-28-10 model on CIFAR-10, encode high-dimensional feature embeddings that capture the data’s underlying structure. To render these high-dimensional features interpretable, we employ a two-step dimensionality reduction process that balances the retention of global structure and the visualization of local relationships.
Setting. The first step in this two-step pipeline is Principal Component Analysis (PCA). By reducing the dimensionality of the latent features to $50$ components, PCA retains the majority of the variance within the data. This step ensures that the global structure of the feature space is preserved, maintaining the relative distances and relationships that reflect the overall organization of the data (Gewers et al., 2021). PCA also acts as a preprocessing step, making the subsequent visualization process more efficient and effective. In the second step, we apply t-Distributed Stochastic Neighbor Embedding (t-SNE) to project the $50$-dimensional PCA-reduced features into a two-dimensional space. Unlike PCA, which emphasizes preserving variance, t-SNE is designed to maintain local relationships among data points (Xia et al., 2021). This combination of PCA and t-SNE allows us to visualize the latent feature space in a manner that is both interpretable and faithful to the original high-dimensional structure.
After deciding on the two-dimensional latent space, we visualize the selected unlabeled data using our proposed schemes: PCS, LCS-KM, and LCS-GMM. To ensure methodological rigor and consistency, we strictly adhere to the procedures outlined in Algorithms 1 and 2 for implementing the data selection techniques. The ratio hyperparameters $\alpha$ and $\beta$ are set to $0.1$ and $1$, respectively. Importantly, these parameters influence the specific distribution of selected points but do not alter the overall interpretability of the visualizations, ensuring that the comparative analysis remains robust by their precise values.
@@ 6.1. Comparative Visualizations
PCS. Figure LABEL:fig:clustering_techniques presents the visualization results, where each color represents the label class predicted by the intermediate model for the corresponding unlabeled point, and the highlighted black dots indicate the data points selected by the respective selection strategy. Our visualization results largely explain why our LCS-KM method consistently achieves the best robustness performance when incorporated in SSAT across various experimental settings. In particular, PCS adopts a purely confidence-based selection strategy, choosing the globally lowest confidence points without considering their distribution within or across classes. Although this direct approach is appealing for its simplicity, it can inadvertently prioritize noisy outliers or ambiguous samples that offer limited benefit for refining the decision boundary. As the number of classes grows and data complexity increases, this scattershot selection tends to yield a more disorganized set of points, as shown in Figure LABEL:fig:PCS_selection. This amplifies PCS’s shortcomings compared to the more structured sampling of LCS-KM or the probabilistic clustering offered by LCS-GMM. Thus, PCS often dedicates training efforts to less informative or highly noisy examples, resulting in weaker robustness improvements.
@@
LCS. LCS-GMM relies on fitting Gaussian distributions to the latent representations. Although this probabilistic approach can effectively capture regions where classes overlap, it inherently presupposes that local data clusters follow (approximately) Gaussian distributions, an assumption that may fail in real-world settings characterized by irregular or noisy data. As a result, LCS-GMM’s contours can become misaligned with the true underlying structure, leading to less accurate localization of essential boundary points (Steinley and Brusco, 2011). A closer look at the selected boundary samples in Figure LABEL:fig:gmm also reveals that the region these points occupy is broader and less tightly focused on the decision boundary, making it harder to capture truly critical examples for training. Due to the concentration, selected data points overlap with each other in the two-dimensional visualization space. In some instances—such as the cyan region in the figure—LCS-GMM fails to cover as many relevant boundary samples as LCS-KM, further underscoring the latter’s advantage in both coverage and precision.
Finally, our best-performing LCS-KM method emphasizes leveraging the latent-space clustering structure to identify critical decision-boundary regions in a more balanced manner. By partitioning the data representations into multiple k-means clusters, LCS-KM systematically locates the points of lowest classification confidence within each cluster, illustrated in Figure LABEL:fig:kmeans, thereby ensuring that the algorithm focuses on diverse boundary-adjacent regions. This cluster-based approach yields a fine-grained analysis of local uncertainties and avoids the risk of overcommitting to isolated, potentially noisy areas in the latent space.
6.2. Hyperparameter Sensitivity
@@
The efficacy of our selection strategies critically relies on several key hyperparameters, including $\beta$ that controls the balance between boundary-adjacent unlabeled data and the remaining data points, the proportion of selected unlabeled data $\alpha$, and the per-batch ratio of extra to original data $\gamma$. We record model robust accuracy by varying each parameter to provide guidelines on how to choose these hyperparameters optimally. Due to page limits, experimental results with varying values of $\gamma$ are deferred to Appendix D.1.
@@
Ratio of Boundary Points $\bm{\beta}$. Concentrating exclusively on points near the decision boundary can result in an excessive number of points close to the boundary and too few farther away, leading to overfitting. To mitigate this, we choose a combination of points, with some near the boundary and others further away. The proportion of these selections is determined by the value of $\beta$. We conduct experiments with different $\beta$ values to assess their impact on robust accuracy. Figure LABEL:fig:betacifar illustrates how robust accuracy varies with $\beta$ on the CIFAR-10 dataset using $20\%$ generated data from DDPM, which is approximately four times the size of the original dataset. In this scenario, the optimal robust accuracy is achieved when $\beta$ is $0.4$, meaning $40\%$ of the selected points are near the boundary, while $60\%$ are farther away. Figure LABEL:fig:betasvhn illustrates how robust accuracy varies with $\beta$ on the SVHN dataset using $10\%$ extra data from the same dataset. Here, due to the smaller amount of external data, selecting points near the boundary results in a lower risk of overfitting. Consequently, the highest robust accuracy is observed when $\beta$ is $0.6$ with $60\%$ of the points near the boundary and $40\%$ farther away. Therefore, the ideal value of $\beta$ depends on the amount of external data added and the ratio of external to original data. When more external data is added, a smaller $\beta$, i.e., fewer points near the boundary, may yield better robust accuracy. Conversely, choosing a larger $\beta$ is favorable with less extra unlabeled data.

### v2

@@ IV-B Computational Benefits of Data Reduction
As suggested by the runtime comparisons demonstrated in Table I, utilizing smaller, carefully selected unlabeled data can significantly reduce the overall training time of SSAT while maintaining competitive robustness. We further visualize the learning curves of SSAT and discuss the computational benefits associated with our proposed methods.
@@
Figure 1 depicts the learning curves of vanilla AT and SSAT algorithms on CIFAR-10 with varying ratios of unlabeled data selected from TinyImages. In our setup, one training epoch corresponds to processing 5050K data points, with $10$ PGD steps for training and $20$ for testing. When no additional data is utilized, the highest robust test accuracy is achieved after approximately $75$ epochs (Figure 1a). Incorporating $10\%$ additional data, whether through random selection or curated approaches, extends the optimal convergence point to approximately $100$ epochs (Figures 1b and 1c). In contrast, leveraging $100\%$ of the unlabeled data delays the achievement of peak accuracy to around $185$ epochs (Figure 1d). These findings suggest that including extra unlabeled data in SSAT increases computational demands. Our selection schemes reduce redundant information in the dataset, allowing SSAT to achieve high robustness performance with fewer training epochs.
@@
Generally speaking, the size of the utilized unlabeled dataset is critical in determining the time required for SSAT to achieve peak robust accuracy. Larger unlabeled datasets usually extend the convergence times, whereas smaller datasets facilitate faster convergence, allowing the model to reach optimal accuracy earlier in training. However, continued training beyond this point often leads to robust overfitting. Early stopping offers a computational advantage in scenarios with limited datasets by preventing overfitting with reduced training durations. Nevertheless, early stopping proves less effective in SSAT when working with large-scale unlabeled datasets, as longer training is necessary to achieve peak performance. To mitigate these challenges, selecting a smaller yet strategically curated subset of data significantly reduces the dataset size, enabling faster convergence and minimizing overall training time. These findings underscore the computational benefits of our proposed data selection schemes, providing a practical solution to the challenges of SSAT with extensive datasets.
(a) Random Selection ($\alpha=10\%$)
(b) LCS-KM ($\alpha=10\%$)
(c) LCS-KM ($\alpha=20\%$)
(d) Full Utilization ($\alpha=100\%$)
Figure 2: Standard and robust accuracy curves of SSAT with labeled data from COVIDGRand unlabeled data from CoronaHack, with different selection schemes and ratios: (a) random selection with $\alpha=10\%$, (b) LCS-KM with $\alpha=10\%$, (c) LCS-KM with $\alpha=20\%$, and (d) all unlabeled data $\alpha=100\%$.
IV-C Guided DDPM Fine-Tuning
We study the performance of our guided DDPM fine-tuning methods as depicted in Algorithm 3 on CIFAR-10. Unlike our previous selection schemes, which required pre-generating the whole dataset, here we generate the same amount of unlabeled data matching the ratio $\alpha$. The data-mixing hyperparameters, namely $\alpha$ and $\beta$, are set consistently with those used in strategic selection. We employ a pretrained DDPM model trained on CIFAR-10 for these guided fine-tuning experiments. Table III presents the results across different $\alpha$ ratios, where our DDPM fine-tuning schemes achieve competitive clean accuracy and robustness performance, matching their selection counterpart. For instance, under the ratio scenario $\alpha=20\%$, leveraging LCG-KM to generate boundary-adjacent data achieves a PGD robust accuracy of $60.3\%$ when integrated in the SSAT framework, which matches the result of $60.2\%$ obtained by LCS-KM. Again, latent clustering with k-means consistently achieves the best robustness compared to the other alternatives. These findings mirror our findings from strategic selection, reinforcing that a small but informative generated set can yield performance comparable to a much larger dataset.
Moreover, we conduct a detailed runtime analysis across different data reduction schemes under the generative setup. In particular, we choose to present the best-performing methods with the optimal hyperparameters, namely LCS-KM and LCG-KM under the scenario of $\alpha=20\%$, and compare the end-to-end training times. We also report the baseline performance of SSAT, where the whole generated dataset ($\alpha=100\%$) and no generated data ($\alpha=0\%$) are used, respectively. The comparison results are summarized in Table II. For clarity, we separately report the computational time of DDPM fine-tuning, data generation, data selection, and the total SSAT training process. Our DDPM generator’s throughput is roughly $71$ images per second, so the approximate data generation times are 200200K images in $46$ minutes, 11M images in $3.9$ hours, and 100100M images in $16.2$ days. As a result, our guided DDPM finetuning (LCG-KM) approach reduces the total runtime from $19.1$ hours to $15.7$ hours when compared to LCS-KM. The improvement will be more pronounced when the total amount of data generated by DDPM is much larger. These results highlight the trade-off between robustness gains and computational cost. While generating millions of synthetic samples can enhance robustness, it is computationally expensive. Our guided fine-tuning approach offers a more resource-efficient alternative, reducing SSAT total runtime from $61.0$ hours to $15.7$ hours while preserving strong robustness.
(a) PCS
(b) LCS-GMM
(c) LCS-KM
Figure 3: Visual comparison of selection techniques on TinyImages dataset in the latent space. Each subplot represents a different method: (a) PCS identifies points with the lowest classification confidence, highlighting areas where the model is most uncertain within the ten-class latent representation, (b) LCS-GMM illustrates probability contours from Gaussian Mixture Models, with selected points emphasizing regions of overlapping probabilities among the ten class clusters, and (c) LCS-KM highlights points selected near decision boundaries across ten classes based on k-means clustering in the latent space.
(a) CIFAR-10
(b) SVHN
(c) CIFAR-10
(d) SVHN
Figure 4: Illustration of SSAT performance using our LCS-KM method on image benchmarks with varying boundary and unlabeled data reduction ratio parameters with respect to: (a) and (c) DDPM-generated CIFAR-10 data with $\alpha=20\%$, (b) and (d) 531531K extra SVHN data with $\alpha=10\%$. For each figure, we vary the considered ratio parameter while keeping all the remaining hyperparameters fixed.
V Additional Analyses
In this section, we present additional analyses to provide further insights into our methods. We first evaluate their performance on a real-world medical application dataset (Section V-A), and then visualize the selected data distribution in the latent space to qualitatively understand their ability in identifying critical data points (Section V-B). Finally, we provide ablation studies on key hyperparameters of our methods. Due to page limit, additional experiments on other hyperparameters, generalizability under different training schemes and perturbations, and the effect of intermediate model quality in low-label regimes, are provided in Appendix D.
V-A Application to Real-World Medical Data
To test the generalization of our findings, we extend our study to a real-world healthcare application. COVIDGR [30] is used as the labeled dataset, and CoronaHack [31] serves as the source of the unlabeled data. Specifically, COVIDGR contains $852$ chest X-ray images, evenly distributed across two classes: $426$ COVID-19 positive (RT-PCR confirmed) and $426$ negative cases. It is then split into $80\%$ training and $20\%$ test sets. CoronaHack comprises $5910$ images with an imbalanced class distribution ($73\%$ positive, $27\%$ negative), split into an $89\%$ training set and an $11\%$ test set. We consider $\ell_{\infty}$ perturbations with $\epsilon=0.1$ and train models on a ResNet-18 architecture using SSAT by applying our LCS-KM strategy to select unlabeled CoronaHack data with varying ratios. Further details of the experimental setup is discussed in Appendix B.
@@ V-A Application to Real-World Medical Data
The results are demonstrated in Figure 2, which confirms the effectiveness of our selection scheme in enhancing robustness while optimizing efficiency. For instance, using LCS-KM to select $10\%$ of the unlabeled data achieves a robust accuracy of $56\%$, compared to $53\%$ when the same proportion is chosen randomly (Figures 2a and 2b). In comparison, using a smaller unlabeled dataset leads to faster convergence, underscoring the benefits of early stopping. Specifically, the best performance of $56\%$ is reached at epoch $36$ when using $10\%$ of the data selected with LCS-KM (Figure 2b). Figure 2c shows that incorporating $20\%$ of the data with LCS-KM improved the accuracy to $57\%$ at epoch $47$. In contrast, training on the entire unlabeled dataset requires $60$ epochs to achieve the peak robust accuracy of $58\%$ (Figure 2d). These findings confirm that using a smaller but strategically selected unlabeled data subset can achieve robustness comparable to that of the whole dataset while significantly reducing computational costs. These results highlight the practical relevance of our data reduction scheme, particularly in medical diagnostics, where labeled data is limited, yet robust model performance is essential.
V-B Comparative Visualizations of Our Methods
To facilitate a deeper understanding, we perform a comparative visualization of the selected unlabeled data on CIFAR-10. We first extract their latent representations from the penultimate layer of a pre-trained WRN-28-10 intermediate model. For visualization, we employ a two-step dimension reduction that balances the retention of global structure and the visualization of local relationships. Specifically, we first perform Principal Component Analysis (PCA) [32]. By reducing the dimensionality of the latent features to $50$, PCA retains the majority of the variance within the data. In the second step, we apply t-Distributed Stochastic Neighbor Embedding (t-SNE) [33] to project the $50$-dimensional PCA-reduced features onto a two-dimensional space. Unlike PCA, which emphasizes preserving variance, t-SNE mainly aims to maintain the local relationships among data points. After projecting onto the two-dimensional space, we visualize the set of unlabeled data returned by our selection schemes under $\alpha=10\%$.
@@ V-B Comparative Visualizations of Our Methods
Figure 3 presents the visualization results, where each color represents the label class predicted by the intermediate model for the corresponding unlabeled point, and the highlighted black dots indicate the data points selected by the respective selection strategy. Our visualization results largely explain why our LCS-KM method consistently achieves the best robustness performance when incorporated in SSAT. In particular, PCS adopts a purely confidence-based selection strategy, choosing the globally lowest confidence points without considering their distribution within or across classes. Although this direct approach is appealing for its simplicity, it can inadvertently prioritize noisy outliers that offer limited benefit for refining the decision boundary. As the number of classes grows and data complexity increases, this scattershot selection tends to yield a more disorganized set of points (Figure 3a).
@@
In comparison, LCS-GMM and LCS-KM yield more structured sampling patterns. While LCS-GMM can roughly capture regions where classes overlap, it assumes that each cluster follows approximately Gaussian distributions, an assumption that may fail in real-world settings. As a result, LCS-GMM’s contours become misaligned with the true underlying structure, leading to less accurate localization of essential boundary points (Figure 3b). A closer look at the selected samples also reveals that the corresponding region is broader and less concentrated on the decision boundary, making it harder to capture truly critical examples for training. In some instances—such as the cyan region in the figure—LCS-GMM fails to cover as many relevant boundary samples as LCS-KM.
Finally, our best-performing LCS-KM method emphasizes the importance of leveraging the latent-space clustering to identify boundary-adjacent data points in a more accurate and balanced manner. By partitioning the latent embeddings into multiple k-means clusters, LCS-KM is able to locate more points with low classification confidence within each cluster (Figure 3c). Compared with LCS-GMM, a notable pattern is that the unlabeled data picked by LCS-KM are aligned more linearly. This coincides with the piecewise linear characteristic of neural networks’ decision boundary, especially for the last layer. We hypothesize that such a local linear property potentially explains why our latent clustering-based approach with k-means always achieves the best performance compared to the other two alternatives, which is worth further investigation.
V-C Hyperparameter Sensitivity
@@
The success of our methods relies on a few key hyperparameters, including $\beta$ that controls the balance between boundary-adjacent unlabeled data and the remaining data points, the proportion of selected unlabeled data $\alpha$, and the per-batch ratio of extra to original data $\gamma$. We record model robust accuracy by varying each parameter to provide guidelines on how to choose these hyperparameters optimally. Due to page limits, the detailed settings of these experiments and the ablations on $\gamma$ are deferred to Appendix C-A.
@@
Ratio of Boundary Data $\bm{\beta}$. Concentrating exclusively on data points near the decision boundary can result in an excessive number of points close to the boundary and too few farther away, leading to overfitting. To mitigate this, we choose a combination of points, with some near the boundary and others further away, where the ratio is determined by $\beta$. Figure 4a illustrates how robust accuracy varies with $\beta$ on CIFAR-10 using $20\%$ generated data from DDPM. In this scenario, the optimal robust accuracy is achieved when $\beta$ is $0.4$, meaning $40\%$ of the selected points are near the boundary, while $60\%$ are farther away. Figure 4b illustrates how robust accuracy varies with $\beta$ on SVHN using $10\%$ extra data from the same dataset. Due to a smaller amount of external data, selecting points near the boundary results in a lower risk of overfitting. As a result, the highest robust accuracy is observed when $\beta$ is $0.6$ with $60\%$ of the points near the boundary and $40\%$ farther away. Therefore, the ideal value of $\beta$ depends on the amount of external data added and the ratio of external to original data. When more external data is added, using a smaller $\beta$ usually leads to better robust accuracy. Conversely, choosing a larger $\beta$ is favorable with less extra unlabeled data.